E=(-4,K) AND F=(K,7) THE SLOPE OF EF IS 2/3, SOLVE FOR K
Answer by stanbon(75887)

E=(-4,K) AND F=(K,7) THE SLOPE OF EF IS 2/3, SOLVE FOR K
----------
slope = (7-K)/(K+4) = 2/3
Cross-multiply to get:
3(7-K) = 2(K+4)
21-3K = 2K+8
5K = 13
K = 13/5
=============
Cheers,
Stan H.